### v3

@@ Abstract
Scene graph generation (SGG) analyzes images to extract meaningful information about objects and their relationships. In the dynamic visual world, it is crucial for AI systems to continuously detect new objects and establish their relationships with existing ones. Recently, numerous studies have focused on continual learning within the domains of object detection and image recognition. However, a limited amount of research focuses on a more challenging continual learning problem in SGG. This increased difficulty arises from the intricate interactions and dynamic relationships among objects, and their associated contexts. Thus, in continual learning, SGG models are often required to expand, modify, retain, and reason scene graphs within the process of adaptive visual scene understanding. To systematically explore Continual Scene Graph Generation (CSEGG), we present a comprehensive benchmark comprising three learning regimes: relationship incremental, scene incremental, and relationship generalization. Moreover, we introduce a “Replays via Analysis by Synthesis" method named RAS. This approach leverages the scene graphs, decomposes and re-composes them to represent different scenes, and replays the synthesized scenes based on these compositional scene graphs. The replayed synthesized scenes act as a means to practice and refine proficiency in SGG in known and unknown environments. Our experimental results not only highlight the challenges of directly combining existing continual learning methods with SGG backbones but also demonstrate the effectiveness of our proposed approach, enhancing CSEGG efficiency while simultaneously preserving privacy and memory usage. All data and source code are publicly available at here.
Keywords:
Scene graph generation Continual learning In-context symbolic replay Long tail distribution Compositional scene graphs
@@ 1 Introduction
The field of continual learning has witnessed significant growth in recent years, with a major focus on tasks such as image classification [37], object detection [58], and visual question answering [26]. However, these endeavors have largely neglected the distinctive complexities associated with CSEGG. Here, we highlight several unique challenges of CSEGG: (1) In contrast to object detection, SGG involves understanding and capturing the relationships between objects, which can be intricate and diverse. Consequently, in CSEGG, conveying the spatial and semantic relationships between objects demands adaptive reasoning from the dynamic scene. (2) SGG introduces a higher level of combinatorial complexity than object detection and image classification because each detected object pair may have multiple potential spatial and functional relationships. Thus, as new objects are introduced to the scenes, the complexity of relationships among all the objects increases significantly in a non-linear fashion. (3) The long-tailed distribution in SGG can be attributed to the inherent characteristics of real-world scenes, where certain objects are more prevalent than others. Consequently, CSEGG requires the computational models to adapt continually to the evolving long-tailed distributions over different scenes. Due to a scarcity of research specifically addressing these challenges of CSEGG, there is a pressing need for specialized investigations and methodologies to enable computational models with the ability of CSEGG.
@@
To address the CSEGG challenges, we present a method called "Replays via Analysis by Synthesis", abbreviated as RAS. RAS employs scene graphs from previous tasks, breaks them down and re-composes them to generate diverse scene structures. These compositional scene graphs are then used for synthesizing scene images for replays. Due to its nature of symbolic replays, RAS doesn’t require the storage of original images, which often carry excessive and redundant details. This also ensures data privacy preservation and data efficiency. Furthermore, by synthesizing scenes using composable scene graphs, RAS maintains the semantic context and structure of previous scenes and also enhances the diversity of scene generation. To prevent biased predictions stemming from long-tailed distributions, we moderate the distribution of replayed scene graphs by balancing tail and head classes. This ensures a uniform sampling of relationships and objects during replays. Extensive experiments underscore the effectiveness of our approach. Network analysis reveals our crucial design choices that can be beneficial for the future development of CSEGG models.
@@ 2 Related Works
Scene Graph Generation Datasets. Visual Phrase [50] stands as one of the earliest datasets in the field of visual phrase recognition and detection. Over time, various large-scale datasets have emerged to tackle the challenges of Scene Graph Generation (SGG) [21, 36, 23, 25, 32, 69, 65, 63, 71, 11, 31, 72]. Among these, the Visual Genome dataset [23] has played a pioneering role by providing rich annotations of objects, attributes, and relationships in images. Despite the significant contributions of these datasets to SGG, none focuses on continual learning in SGG. As the preliminary efforts towards CSEGG, we re-structure the Visual Genome dataset [23] and establish a novel and comprehensive CSEGG benchmark, where AI models are deployed to dynamic scenes where new objects and new relationships are introduced.
@@
Continual Learning Methods. Existing continual learning works can be categorized into several approaches. (1) Regularization-based methods [22, 9, 70, 2, 4] aim to mitigate catastrophic forgetting by employing regularization techniques in the parameter space. (2) Dynamic architecture-based approaches[57, 67, 20, 41] adapt the model’s architecture dynamically to accommodate new tasks without interfering with the existing ones. (3) Replay-based methods [48, 10, 47, 56, 44, 7] utilize a memory buffer to store and replay past data during training, enabling the model to revisit and learn from previously seen examples, thereby reducing forgetting. The special variants of these methods include generative replay methods, such as [52, 62, 66, 43], where synthetic data is generated and replayed. Although these generative replay methods, as well as other continual learning methods, have been extensively studied in image classification [8, 61, 37] and object detection[59, 51, 39], few works focus on the challenges in CSEGG, such as adaptive reasoning from the dynamic scenes, the evolving long-tailed distribution across scenes, and the combinatorial complexity involving objects and their multiple relationships. In this work, we introduce a continual learning method, abbreviated as RAS (Replays via Analysis by Synthesis). To address the distinct challenges of CSEGG, RAS involves creating in-context synthetic scene images based on re-composable scene graphs from previous tasks to reinforce continual learning. The components in RAS facilitate memory-efficient training and preserve privacy while maintaining the scene diversity and scene context for SGG in dynamic environments.
Scenarios #Tasks #Objs #Rels
@@
SGG
backbones
CL
baselines
S1 5 All tasks: 150 Each task: 10 F, R, mF, mR, FWT, BWT, R_bbox, Gen_R Transformer based (SGTR) CNN based (TCNN) Joint, Naive, Replays[48], EWC[22], PackNet[38]
@@
S3 4 Each task: 30 All tasks: 35
Table 1: Overview of three CSEGG learning scenarios. This table summarizes the three learning scenarios in CSEGG, including the number of tasks, the number of object (#Objs) and relationship (#Rels) classes, the evaluation metrics, the SGG-Backbones used, and the continual learning (CL) baselines. See Sec. 3 for details.
@@ 3.1 Learning Scenarios
Figure 2: Three learning scenarios are introduced. From left to right, they are S1. relationship (Rel.) incremental learning (Incre.); S2. scene incremental learning; and S3. relationship generalization (Rel. Gen.) in Object Incre.. In S1 and S2, example triplet labels in the training (solid line) and test sets (dash line) from each task are presented. The training and test sets from the same task are color-coded. The new objects or relationships in each task are bold and underlined. In S3, one single test set (dashed gray box) is used for benchmarking the relationship generalization ability of object incre. learning models across all the tasks.
@@
Scenario 1 (S1): Relationship Incremental Learning. To uncover contextual information and go beyond studies of object detection and recognition, we introduce this scenario consisting of 5 tasks where 10 new relationship classes are incrementally added in every task (Fig. 2, left; Tab. 1). All object classes and their locations are made known to all CSEGG models over all the tasks. This scenario resembles a human learning scenario where a parent gradually teaches a baby to recognize new relationships among all objects in the same room, focusing on one new relationship at a time during continual learning. This scenario also has implications in medical imaging where identical cell types may form new relationships with nearby cells depending on the context. See Sec. 0.A.1.1 for details.
@@
Scenario 2 (S2): Scene Incremental Learning. To simulate the real-world cases when there are demands for detecting new objects and new relationships from old to new scenes, we introduce this scenario where new objects and new relationships are incrementally introduced over tasks (Fig. 2, middle; Tab. 1). There are 2 tasks in total with the first task containing 100 object classes and 40 relationship classes with 25 more object classes and 5 more relationship classes in the second task. This aligns with the real-world use cases where common objects and relationships are learned in the first scene, and incremental learning in the second scene only happens on less frequent relationships and objects. See Sec. 0.A.1.2 for details.
@@
Scenario 3 (S3): Relationship Generalization. Humans have no problem at all recognizing the relationships of unknown objects with other nearby objects. This scenario is designed to investigate the relationship generalization ability of CSEGG models. This capability is essential for real-world implications, such as in robotic navigation where it often encounters unknown objects and requires classifying their relationships. In total, there are four tasks, each introducing an incremental addition of 30 new object classes. All relationship classes are made known to all CSEGG models over all the tasks (Fig. 2, right; Tab. 1). Different from scenarios S1 and S2, a standalone generalization test set is curated, where the objects are unknown but the relationship classes among these unknown objects are common to the training set of every task. The CSEGG models trained after every task are tested on this standalone generalization test set to predict relationships among the unknown objects. See Sec. 0.A.1.3 for details.
@@
Data sampling and distributions. To allocate data for every task of each scenario, we perform the following sampling strategies. In S1 and S3 above, either object or relationship classes are randomly sampled from the Visual Genome dataset and incrementally added to every task. Due to the inherent characteristics of real-world scenes, the long-tailed class distribution is present in S1 and S3. However, in S2, only tail classes are sampled and added in subsequent tasks.
@@ 3.2 Competitive CSEGG Baselines
Due to the scarcity of CSEGG works, we contribute a diverse set of competitive CSEGG baselines and implement them on our own. Each CSEGG baseline requires three components: a backbone model for scene graph generation (SGG), a continual learning (CL) method to prevent the SGG model from forgetting, and an optional data sampling technique to deal with imbalanced data at every task for training SGG models. Next, we introduce the 2 SGG backbones, the 5 continual learning methods, and the 5 optional data sampling techniques. See Sec. 0.A.2 for implementation and training details of CSEGG baselines.
SGG Backbones. We use the two state-of-the-art backbones: (1) one-stage Scene graph Generation TRansformer (SGTR) [28] and (2) the traditional Two-stage SGG model (TCNN) [63]. Briefly, SGTR (Fig. S2(a)) uses a transformer-based architecture for image feature extraction and fusion. During training, [28] formulates SGG as a bipartite graph construction and matching problem. In contrast, TCNN detects objects with Faster-RCNN[17] backbone and infers their relationships separately via Iterative message passing [63]. We use public source codes from [28] and [59] with default hyperparameters.
Continual Learning Methods. We include the following continual learning methods (Fig. S2(b)): (1) Naive (lower bound) is trained on each task in sequence without any measures to prevent catastrophic forgetting. (2) EWC[22] is a weight-regularization method, where the weights of the network are regularized in the parameter space, based on their “importance" to the previous tasks. (3) PackNet[38] is a parameter-isolation method, iteratively pruning the network parameters after every task, so that it can sequentially pack multiple tasks within one network. (4) Replay@M[48] includes a memory buffer with the capacity of storing $M$ percentages of images in the entire dataset as well as their corresponding ground truth object and predicate notations depending on the task at each learning scenario. We vary $M=$ 10%, 20%, and 100%. (5) Joint Training is an upper bound where the SGG model is trained on the entire CSEGG dataset.
@@ 3.3 Evaluation Metrics
In scenario S3, we evaluate all CSEGG methods in the standalone generalization test set, shared over all the tasks. To benchmark generalization abilities in unknown object localization and relationship classification among these unknown objects, we introduce two evaluation metrics: Gen Rbbox@K and Gen R@K. As the CSEGG models have never been taught to classify unknown objects, we discard the class labels of the bounding boxes and only evaluate the predicted bounding box locations with Gen Rbbox@K. To evaluate whether the predicted bounding box location is correct, we apply a hard threshold of Intersection over Union (IoU) between the predicted bounding box locations and the ground truth. Any predicted bounding boxes with their IoU values above the hard threshold are deemed to be correct. We vary IoU thresholds from 0.3, 0.5, to 0.7. To assess whether the CSEGG model generalizes to detect known relationships over unknown objects, we evaluate the recall Gen R@K of the predicted relationships $r_{k}$ only on correctly predicted bounding boxes. See Sec. 0.A.3 for detailed definitions of these metrics. All experimental results are averaged over three runs.
@@ 4 Replays via Analysis by Synthesis (RAS)
To address the complexities in CSEGG, we introduce our “Replays via Analysis by Synthesis" method, dubbed RAS. Our RAS belongs to the group of generative replay methods for continual learning. However, different from the existing generative replay methods [52, 14, 15], our RAS leverages symbolic replays with state-of-the-art diffusion models. Moreover, rather than generating any random images for replays, our RAS is capable of generating in-context images following semantic rules, such as object co-occurrences. The schematic of our RAS is presented in Fig. 3. Next, we focus on how RAS generates images and predicts SGG annotations on these images for replays.
@@
To generate images $I^{\prime}_{j}$ for replays, our RAS feeds text prompts, which are formed by a set of chosen triplet labels and describe the diverse in-context scenes, into the state-of-the-art Stable Diffusion model [49]. The small object on the table is most likely to be a spoon rather than an elephant. Context plays important roles in visual perceptions [5]. To generate text prompts describing context-congruent scenes, we employ a context checker. First, the context checker uses the pre-trained large language model BERT [13] to extract embeddings for each triplet label in $U_{t}$. As BERT has been pre-trained on a large corpus of text data, it learns to capture context-relevant representations of words. Next, hierarchical clustering is performed on these embeddings using the agglomerative clustering algorithm [40]. This ensures that each cluster contains only embeddings that are semantically close. The threshold for the agglomerative clustering algorithm is set to 0.6. As real-world images often contain complex scenes involving multiple triplets, we select any cluster with more than 3 triplet labels to create a text prompt for the image generation model.
@@
Figure 4: Visualizations of example generated images given the text prompt from our RAS. Each row presents two examples of generated images. Given the input text prompt on the left, the generated image is displayed on the right. These visualizations provide qualitative validations into the capability of RAS to produce diverse and contextually relevant images based on our designed text prompts.
We provide the visualization examples of some synthesized images along with the corresponding text prompts in Fig. 4. From these examples, we found that the composed text prompts and the synthesized images are often of high quality and contextual coherence, thanks to the Stable Diffusion Model [49].
Scene Graph Prediction on Synthesized Images. During replays, to train the model $M_{t+1}$ on $I^{\prime}_{j}$, we also need to predict their corresponding scene graph notations $G^{\prime}_{j}$ on $I^{\prime}_{j}$. As the frozen model snapshot $M_{t}$ at the end of task $t$ carries prior knowledge for SGG from the previous tasks, we use it to predict notations $G^{\prime}_{j}$ on $I^{\prime}_{j}$. These $G^{\prime}_{j}$ comprises object nodes $O^{\prime}_{j}$ with their respective classes $c^{\prime}_{j}$, along with object locations $b^{\prime}_{j}$. Additionally, it includes corresponding relationship nodes $R^{\prime}_{j}$ formed by triplets <$o^{\prime}_{s}$, $p^{\prime}_{k}$, $o^{\prime}_{j}$> representing subject, predicate, and object nodes, respectively. These generated notations $G^{\prime}_{j}$, along with $I^{\prime}_{j}$, serve to construct the exemplars, used for replays.

### v4

@@ Abstract
Scene graph generation (SGG) analyzes images to extract meaningful information about objects and their relationships. In the dynamic visual world, it is crucial for AI systems to continuously detect new objects and establish their relationships with existing ones. Recently, numerous studies have focused on continual learning within the domains of object detection and image recognition. However, a limited amount of research focuses on a more challenging continual learning problem in SGG. This increased difficulty arises from the intricate interactions and dynamic relationships among objects, and their associated contexts. Thus, in continual learning, SGG models are often required to expand, modify, retain, and reason scene graphs within the process of adaptive visual scene understanding. To systematically explore Continual Scene Graph Generation (CSEGG), we present a comprehensive benchmark comprising three learning regimes: relationship incremental, scene incremental, and relationship generalization. Moreover, we introduce a “Replays via Analysis by Synthesis" method named RAS. This approach leverages the scene graphs, decomposes and re-composes them to represent different scenes, and replays the synthesized scenes based on these compositional scene graphs. The replayed synthesized scenes act as a means to practice and refine proficiency in SGG in known and unknown environments. Our experimental results not only highlight the challenges of directly combining existing continual learning methods with SGG backbones but also demonstrate the effectiveness of our proposed approach, enhancing CSEGG efficiency while simultaneously preserving privacy and memory usage. All data and source code are publicly available here.
@@ 1 Introduction
The field of continual learning has witnessed significant growth in recent years, with a major focus on tasks such as image classification [43], object detection [67], and visual question answering [28]. However, these endeavors have largely neglected the distinctive complexities associated with CSEGG. Here, we highlight several unique challenges of CSEGG: (1) In contrast to object detection, SGG involves understanding and capturing the relationships between objects, which can be intricate and diverse. Consequently, in CSEGG, conveying the spatial and semantic relationships between objects demands adaptive reasoning from the dynamic scene. (2) SGG introduces a higher level of combinatorial complexity than object detection and image classification because each detected object pair may have multiple potential spatial and functional relationships. Thus, as new objects are introduced to the scenes, the complexity of relationships among all the objects increases significantly in a non-linear fashion. (3) The long-tailed distribution in both objects and relationships in SGG can be attributed to the inherent characteristics of real-world scenes, where certain objects are more prevalent than others. Consequently, CSEGG requires the computational models to adapt continually to the evolving long-tailed distributions over different scenes. Due to a scarcity of research specifically addressing these challenges of CSEGG, there is a pressing need for specialized investigations and methodologies to enable computational models with the ability of CSEGG.
@@
To address the CSEGG challenges, we present a method called "Replays via Analysis by Synthesis", abbreviated as RAS. RAS employs scene graphs from previous tasks, breaks them down and re-composes them to generate diverse scene structures. These compositional scene graphs are then used for synthesizing scene images for replays. Due to its nature of symbolic replays, RAS does not require the storage of original images, which often carry excessive and redundant details. This also ensures data privacy preservation and data efficiency. Furthermore, by synthesizing scenes using composable scene graphs, RAS maintains the semantic context and structure of previous scenes and also enhances the diversity of scene generation. To prevent biased predictions stemming from long-tailed distributions, we moderate the distribution of replayed scene graphs by balancing tail and head classes. This ensures a uniform sampling of relationships and objects during replays. Extensive experiments underscore the effectiveness of our approach. Network analysis reveals our crucial design choices that can be beneficial for the future development of CSEGG models.
@@ 2 Related Works
Scene Graph Generation Datasets. Visual Phrase [59] stands as one of the earliest datasets in the field of visual phrase recognition and detection. Over time, various large-scale datasets have emerged to tackle the challenges of Scene Graph Generation (SGG) on static images [23, 42, 25, 27, 37, 78, 74, 72, 80, 12, 35, 81]. Subsequent works further extend the SGG to dynamic videos [22, 50, 56]. Despite the significant contributions of these datasets to SGG, none focuses on continual learning in SGG. As the preliminary efforts towards CSEGG, we start with fundamental and straightforward settings of SGG on static images. Among all the SGG datasets on static images, the Visual Genome dataset [25] has played a pioneering role by providing rich annotations of objects, attributes, and relationships in images. Thus, we re-structure the Visual Genome dataset [25] and establish a novel and comprehensive CSEGG benchmark, where AI models are deployed to dynamic scenes where new objects and new relationships are introduced.
@@
S. #Tasks #Objs #Rels
@@
SGG Backbone CL base. Kn. Unk.
S1 5 150 (All) 10 per task F, R, mF, mR, FWT, BWT, Gen Rbbox, Gen R Transformer based (SGTR) CNN based (IMP) Joint, Naive, Replay M%, EWC, PackNet, RAS_GT
Objs
(bbox, labels)
Rels
@@
None
Rels and Objs
(bbox, labels)
S3 4 30 per task 35 per task
Rels
Objs
(bbox, labels)
Table 1: Overview of three CSEGG learning scenarios. This table summarizes the three learning scenarios (Column 1) in CSEGG, including the number of tasks, the number of object (#Objs) and relationship (#Rels) classes, the evaluation metrics, the SGG-Backbones used, and the continual learning (CL) baselines. The Kn. and Unk. columns provide information regarding what is known to the CSEGG models during training in that scenario and what is being incrementally learned by the models. Unknown information is being incrementally learned by the models. See Sec. 3 for details.
Continual Learning Methods. Existing continual learning works can be categorized into several approaches. (1) Regularization-based methods [24, 9, 79, 2, 4] aim to mitigate catastrophic forgetting by employing regularization techniques in the parameter space. (2) Dynamic architecture-based approaches[66, 76, 21, 47] adapt the model’s architecture dynamically to accommodate new tasks without interfering with the existing ones. (3) Replay-based methods [57, 10, 55, 65, 52, 7] utilize a memory buffer to store and replay past data during training, enabling the model to revisit and learn from previously seen examples, thereby reducing forgetting. The special variants of these methods include generative replay methods, such as [61, 71, 75, 51], where synthetic data is generated and replayed. Although these generative replay methods, as well as other continual learning methods, have been extensively studied in image classification [8, 70, 43] and object detection[68, 60, 45], few works focus on the challenges in CSEGG, such as adaptive reasoning from the dynamic scenes, the evolving long-tailed distribution across scenes, and the combinatorial complexity involving objects and their multiple relationships. In this work, we introduce a continual learning method, abbreviated as RAS (Replays via Analysis by Synthesis). To address the distinct challenges of CSEGG, RAS involves creating in-context synthetic scene images based on re-composable scene graphs from previous tasks to reinforce continual learning. The components in RAS facilitate memory-efficient training and preserve privacy while maintaining the scene diversity and scene context for SGG in dynamic environments. With the rise of pretrained vision-language models (VLMs) [49, 31, 83, 40], various SGG methods [11, 30] have been proposed to tackle open-vocabulary and zero-shot SGG challenges. However, these settings differ fundamentally from CSEGG, where we aim to simulate scenarios where the model encounters novel predicates or objects, unseen by any models, including LLMs or multi-modal models. Using pre-learned information from LLMs or frozen encoders conflicts with the continual learning setting we address in CSEGG.
@@ 3.1 Learning Scenarios
Figure 2: Three learning scenarios are introduced. From left to right, they are S1. relationship (Rel.) incremental learning (Incre.); S2. scene incremental learning; and S3. relationship generalization (Rel. Gen.) in Object Incre.. In S1 and S2, example triplet labels in the training (solid line) and test sets (dotted line) from each task are presented. The training and test sets from the same task are color-coded. Blue color indicates task 1 and orange color indicates task 2. The new objects or relationships in each task are bold and underlined. In S3, one single test set (dotted gray box) is used for benchmarking the relationship generalization of object incre. learning models across all the tasks.
@@
Scenario 1 (S1): Relationship Incremental Learning. To uncover contextual information and go beyond studies of object detection and recognition, we introduce this scenario consisting of 5 tasks where 10 new relationship classes are incrementally added in every task (Fig. 2, left; Fig. S1; Tab. 1). All object classes and their locations are made known to all CSEGG models over all the tasks. This scenario resembles a human learning scenario where a parent gradually teaches a baby to recognize new relationships among all objects in the same room, focusing on one new relationship at a time during continual learning. This scenario also has implications in medical imaging where identical cell types may form new relationships with nearby cells depending on the context (Sec. A.1.1).
@@
Scenario 2 (S2): Scene Incremental Learning. To simulate the real-world cases when there are demands for detecting new objects and new relationships from old to new scenes, we introduce this scenario where new objects and new relationships are incrementally introduced over tasks (Fig. 2, middle; Fig. S1; Tab. 1). There are 2 tasks in total with the first task containing 100 object classes and 40 relationship classes with 25 more object classes and 5 more relationship classes in the second task. This aligns with the real-world use cases where common objects and relationships are learned in the first scene, and incremental learning in the second scene only happens on less frequent relationships and objects. See Sec. A.1.2 for details.
@@
Scenario 3 (S3): Relationship Generalization. Humans have no problem at all recognizing the relationships of unknown objects with other nearby objects. This scenario is designed to investigate the relationship generalization ability of CSEGG models. This capability is essential for real-world implications, such as in robotic navigation where it often encounters unknown objects and requires classifying their relationships. In total, there are four tasks, each introducing an incremental addition of 30 new object classes. All relationship classes are made known to all CSEGG models over all the tasks (Fig. 2, right; Fig. S1; Tab. 1). Different from scenarios S1 and S2, a standalone generalization test set is curated, where the objects are unknown but the relationship classes among these unknown objects are common to the training set of every task. The CSEGG models trained after every task are tested on this standalone generalization test set to predict relationships among the unknown objects. See Sec. A.1.3 for details.
@@
Data sampling and distributions. To allocate data for every task of each scenario, we perform the following sampling strategies. In S1 and S3 above, either object or relationship classes are randomly sampled from the Visual Genome dataset and incrementally added to every task. Due to the inherent characteristics of real-world scenes, the long-tailed class distribution is present in S1 and S3. However, in S2, only tail classes are sampled and added in subsequent tasks. The number of tasks in each scenario is experimentally determined to optimize the training data configuration, ensuring sufficient training samples in each task while maximizing the number of tasks. For detailed statistics, see Fig. S2 and Sec. A.2.
@@ 3.2 Competitive CSEGG Baselines
Due to the scarcity of CSEGG works, we contribute a diverse set of competitive CSEGG baselines and implement them on our own. Each CSEGG baseline requires three components: a backbone model for scene graph generation (SGG), a continual learning (CL) method to prevent the SGG model from forgetting, and an optional data sampling technique to deal with imbalanced data at every task for training SGG models. Next, we introduce the 2 SGG backbones, the 5 continual learning methods, and the 5 optional data sampling techniques. See Sec. A.3 for implementation and training details of CSEGG baselines.
SGG Backbones. We use the two state-of-the-art backbones: (1) one-stage Scene graph Generation TRansformer (SGTR) [32] and (2) the traditional Two-stage SGG model (TCNN) [72]. Briefly, SGTR (Fig. S3 left) uses a transformer-based architecture for image feature extraction and fusion. During training, [32] formulates SGG as a bipartite graph construction and matching problem. In contrast, TCNN detects objects with Faster-RCNN[18] backbone and infers their relationships separately via Iterative message passing [72]. We use implementations from [32] and [68] with default hyperparameters.
Baselines. We include the following continual learning methods (Fig. S3 right): (1) Naive (lower bound) is trained on each task in sequence without any measures to prevent catastrophic forgetting. (2) EWC[24] is a weight-regularization method, where the weights of the network are regularized in the parameter space, based on their “importance" to the previous tasks. (3) PackNet[44] is a parameter-isolation method, iteratively pruning the network parameters after every task, so that it can sequentially pack multiple tasks within one network. (4) Replay@M[57] includes a memory buffer with the capacity of storing $M$ percentages of images in the entire dataset as well as their corresponding ground truth object and predicate notations depending on the task at each learning scenario. We vary $M=$ 10%, 20%, and 100%. (5) Joint Training is an upper bound where the SGG model is trained on the entire CSEGG dataset. (6) RAS_GT is a baseline in which we use the ground truth scene graph labels from each task to create replay buffers using an image generation model explained in detail in Sec. 4. See Fig. S4 for schematics of CSEGG baselines. We provide mathematical formulations of these baselines in Sec. A.4.
@@ 3.3 Evaluation Metrics
In scenario S3, we evaluate all CSEGG methods in the standalone generalization test set, shared over all the tasks. To benchmark generalization abilities in unknown object localization and relationship classification among these unknown objects, we introduce two evaluation metrics: Gen Rbbox@K and Gen R@K. As the CSEGG models have never been taught to classify unknown objects, we discard the class labels of the bounding boxes and only evaluate the predicted bounding box locations with Gen Rbbox@K. To evaluate whether the predicted bounding box location is correct, we apply a hard threshold of Intersection over Union (IoU) between the predicted bounding box locations and the ground truth. Any predicted bounding boxes with their IoU values above the hard threshold are deemed to be correct. We vary IoU thresholds from 0.3, 0.5, to 0.7.
To assess whether the CSEGG model generalizes to detect known relationships over unknown objects, we evaluate the recall Gen R@K of the predicted relationships $r_{k}$ only on correctly predicted bounding boxes. See Sec. A.5 for details. All results are averaged over 3 runs.
@@ 4 Replays via Analysis by Synthesis (RAS)
To address the complexities in CSEGG, we introduce our “Replays via Analysis by Synthesis" method, dubbed RAS. Our RAS belongs to the group of generative replay methods for continual learning. We create an exemplar set $E_{t}$ for replays. At task $t$, we jointly train the scene graph generation model $M_{t}$ on $E_{t}$ and the current training dataset $D_{t}$. However, different from the existing generative replay methods [61, 15, 16], our RAS leverages symbolic replays with state-of-the-art diffusion models. Moreover, rather than generating any random images for replays, our RAS is capable of generating in-context images following semantic rules, such as object co-occurrences. The schematic of our RAS is presented in Fig. 3. Next, we focus on how RAS creates $E_{t}$ containing the generated images and their SGG annotations on these images for replays.
@@
To generate images $I^{\prime}_{j}$ for replays, our RAS feeds text prompts, which are formed by a set of chosen triplet labels and describe the diverse in-context scenes, into the state-of-the-art Stable Diffusion model [58]. As previous works suggests context plays important roles in visual perceptions [5]. To generate text prompts describing context-congruent scenes, we employ a context checker. First, the context checker uses the pre-trained large language model BERT [14] to extract embeddings for each triplet label in $U_{t}$. As BERT has been pre-trained on a large corpus of text data, it learns to capture context-relevant representations of words. Next, hierarchical clustering is performed on these embeddings using the agglomerative clustering algorithm [46]. This ensures that each cluster contains only embeddings that are semantically close. The threshold for the agglomerative clustering algorithm is set to 0.6. As real-world images often contain complex scenes involving multiple triplets, we select any cluster with more than 3 triplet labels to create a text prompt for image generation.
@@
We provide the visualization examples of some synthesized images along with the corresponding text prompts in Fig. S5. From these examples, we found that the composed text prompts and the synthesized images are often of high quality and contextual coherence.
Scene Graph Prediction on Synthesized Images. During replays, to train the model $M_{t+1}$ on $I^{\prime}_{j}$, we also need to predict their corresponding scene graph notations $G^{\prime}_{j}$ on $I^{\prime}_{j}$. As the frozen model snapshot $M_{t}$ at the end of task $t$ carries prior knowledge for SGG from the previous tasks, we use it to predict notations $G^{\prime}_{j}$ on $I^{\prime}_{j}$. These $G^{\prime}_{j}$ comprises object nodes $O^{\prime}_{j}$ with their respective classes $c^{\prime}_{j}$, along with object locations $b^{\prime}_{j}$. Additionally, it includes corresponding relationship nodes $R^{\prime}_{j}$ formed by triplets <$o^{\prime}_{s}$, $p^{\prime}_{k}$, $o^{\prime}_{j}$> representing subject, predicate, and object nodes, respectively. These generated notations $G^{\prime}_{j}$, along with $I^{\prime}_{j}$, serve to construct the exemplars $E_{t}$, used for replays.